# Phase error rate estimation with basis-efficiency mismatch for decoy-state BB84

### v1

@@ VI.4 Bounds on zero and one photon statistics
The lower bound on the zero-photon component is given by [28, Eq. 2]
@@
The lower bound on the one-photon component is given by [28, Eq. 3]
@@
The upper bound on the one-photon component is given by [28, Eq. 4]
@@
Since $\Pr(\mathcal{S}_{\text{constraints}})\geq 1-9\varepsilon_{\text{AT-d}}^{2}$, and Eqs. 59, 60 and 61 follow from the expressions in $\mathcal{S}_{\text{constraints}}$. We have
@@
VI.5 Variable-length security proof
Having proved Eq. 48, we will now use it to prove the variable-length security of the decoy-state BB84 protocol (Theorem 3). Similar to the proof of Theorem 1, we will use $\kappa(\widetilde{\Omega})$, to denote the probability of our computed bounds failing conditioned on the event $\widetilde{\Omega}$. We will see that the events that we will need to condition on are given by $\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})}$. Even though we do not have access to the value $n_{K,{1}}$, we will see that this is the “right” event to condition on. We do not generate key from the zero-photon component $n_{K,{0}}$, since it has little impact on key rates for typical use cases. To see how one may do so, see Remark 17. In the remainder of this section, we identify $\widetilde{\Omega}=\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})}$ for brevity. As in Section III, we define
which when combined with Eq. 48 implies
Thus, the average probability of either of our bounds failing (for any attack undertaken by Eve) is small. We now state and prove the following theorem concerning the variable-length security of decoy-state protocol.
Theorem 3 ( Variable-length security of decoy-state BB84).
Suppose Eq. 48 is satisfied and let $\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ be a function that determines the number of bits used for error-correction in the QKD protocol. Define
$l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ :-$\coloneq\max\Bigg{(}0,\mathcal{B}_{1}\left(n_{K,\mu_{\vec{k}}}\right)\left(1-h\left(\mathcal{B}_{e}\left(e^{\text{obs}}_{X,\mu_{\vec{k}}},n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}}\right)\right)\right)-\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}),:$ (65)
$-2\log(1/2\varepsilon_{\text{PA}})-\log(2/\varepsilon_{\text{EV}})\Bigg{)}$
@@ VI.5 Variable-length security proof
Proof.
Again, the proof is similar to that of Theorem 1 with some differences, and consists of two parts. In the first part, we use Eq. 48 along with the entropic uncertainty relations to bound the smooth min-entropy of the raw key register. Here the main difference is the use of entropic chain rules to isolate the single-photon component of the raw key, before applying the EUR statement. In the second part, we use the obtained bound to prove the variable-length security statement. Here the main difference is due to to the presence of the $n_{K,{1}}$ in the event $\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{0}},n_{K,{1}})}=\widetilde{\Omega}$, which is handled differently. Let us first focus on the event $\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{0}},n_{K,{1}})}=\widetilde{\Omega}$.
where we used [35, Lemma 6.7] to isolate the single-photon contribution to the key in the first inequality, and the entropic uncertainty relations [1] followed by the same series of steps as in the proof of Theorem 1 to replace the smooth max entropy term with the bound on the phase error rate. This is the required bound on the smooth min entropy of the raw key.
Remark 17.
$=\Delta_{1}+\Delta_{2}$
Here, we used the leftover-hashing lemma [2, Proposition 9] on the sub-normalized state $(\rho_{|\widetilde{\Omega}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality, and [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [35, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 66 for the fourth inequality. The fifth equality follows by replacing the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 65. We use the concavity of the square root function and Jensen’s inequality for the sixth inequality to pull the sum over the events and the probability inside the square root, while the seventh equality follows simply from the definition of conditional probabilities and $\kappa$ (Eq. 63).
Combining our bounds on $\Delta_{1}$ and $\Delta_{2}$ from Eqs. 70 and 69, we obtain
$\leq\varepsilon_{\text{PA}}+2\varepsilon_{\text{AT}}.$
@@
For instance, one way to avoid this problem is to argue that Bob can compare his raw key before and after error-correction to calculate the number of errors in the key generation rounds (assuming error-correction succeeded). This can indeed be made rigorous by arguing that if error-correction fails, the protocol aborts with high probability anyway (due to error-verification). However an additional issue remains. For variable-length protocols, Bob must announce either the number of errors he observes, or the length of key he wishes to produce, to Alice. This additional announcement leaks information to Eve which must be accounted for. Assuming that Bob announces the final output key length, this would reduce the key length by an additional $\log(l_{\text{max}}+1)$ where $l_{\text{max}}$ denotes the maximum allowed output key length ($(l_{\text{max}}+1)$ since abort is modelled as key length of 00). This observation is missing in Ref. [50].
@@
Note that this problem is avoided by this work, and by [28], since the length of output key is a function only of the public announcements during Step 3 of the protocol (Section VI.1).
VII Results
@@
We will now apply our results to a decoy-state BB84 protocol with realistic detectors. To do so, we start by specifying the canonical model for our detectors in Section VII.1. We will then specify a recipe for using this work to compute key rates in Section VII.2. Finally, we will plot the key rate we obtain in Section VII.3.
VII.1 Detector Model
@@
Let $\eta_{b_{i}},d_{b_{i}}$ denote the efficiency and dark count rate of Bob’s POVM corresponding to basis $b$, and bit $i$. We first define Bob’s double click POVM for basis $b\in\{Z,X\}$ to be $\Gamma^{(B)}_{(b,\text{dc})}=\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{0}})(1-\eta_{b_{0}})^{N_{0}})(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ , where |$\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ is the state with $N_{0}$ photons in the mode 1, and $N_{1}$ photons in mode 2, where the modes are defined with respect to basis $b$. For example, for polarisation-encoded BB84, |$\outerproduct{2,1}{2,1}_{Z}$ would signify the state with $2$ horizontally-polarised photons and $1$ vertically polarised photon. Recall that double clicks are mapped to single clicks randomly in our protocol. Thus, we can write Bob’s POVM elements as
@@
$\Gamma^{(B)}_{(b,1)}$ $=(1-d_{b_{0}})\sum_{N_{0},N_{1}=0}^{\infty}(1-\eta_{b_{0}})^{N_{0}}(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}+\frac{1}{2}\Gamma^{(B)}_{(b,\text{dc})},$
Decoy methods allow us to restrict out attention to rounds where Alice sent single photons. Thus her Hilbert space is qubit while Bob holds two optical modes. The joint Alice-Bob for the basis $b$ can be constructed via Eqs. 2, 3 and 72 and is given by
@@
VII.2 Recipe for computing key rates in the presence of basis-efficiency mismatch
In this subsection, we provide straightforward instructions for using the results of this work to compute key rates for decoy-state BB84 in the presence of basis-efficiency mismatch (see the end of this subsection for a pointer to the exact expressions). We will start by explaining the computation of (upper bounds on) $\delta_{1},\delta_{2}$ for a given model. To do so, one has to break up the measurement process implemented by Alice and Bob into multiple steps via multiple uses of Lemma 1. This is done as follows:
  1. 1.
    Start with POVM $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ (from Eq. 72) which describe Alice and Bob measuring in the $(b_{A},b_{B})$ basis, and obtaining a conclusive error, a conclusive no-error, and an inconclusive outcome respectively.
  2. 2.
    This POVM measurement can be divided into two steps. The first step is implemented using $\{\tilde{F}_{(b_{A},b_{B}),(\text{con})},\tilde{F}_{(b_{A},b_{B}),(\bot)}\}$ and decides whether the outcome is conclusive or inconclusive. The conclusive outcomes are further measured using $\{G^{\text{con}}_{(b),(\neq)},G^{\text{con}}_{(b),(=)}\}$. These POVM elements are given by
    $\tilde{F}_{(b_{A},b_{B}),(\text{con})}$ $=\Gamma_{(b_{A},b_{B}),(\neq)}+\Gamma_{(b_{A},b_{B}),(=)}=\mathrm{I}_{A}\otimes(\mathrm{I}_{B}-\Gamma^{(B)}_{(b_{B},\bot)})$ (75)
    $\tilde{F}_{(b_{A},b_{B}),(\bot)}$ $=\Gamma_{(b_{A},b_{B}),(\bot)}=\mathrm{I}_{A}\otimes\Gamma^{(B)}_{(b_{B},\bot)}$
    $G^{\text{con}}_{(b),(\neq)}$ $=\sqrt{\tilde{F}_{(b),(\text{con})}}^{+}\Gamma_{(b),(\neq)}\sqrt{\tilde{F}_{(b),(\text{con})}}^{+}$
    $G^{\text{con}}_{(b),(=)}$ $=\sqrt{\tilde{F}_{(b),(\text{con})}}^{+}\Gamma_{(b),(=)}\sqrt{\tilde{F}_{(b),(\text{con})}}^{+}+\mathrm{I}-\mathrm{\Pi}_{\tilde{F}_{(b),(\text{con})}}\qquad\qquad(=\mathrm{I}-G^{\text{con}}_{(b),(\neq)})$
    where $b$ in the above expressions denotes the basis used by both Alice and Bob, and $\mathrm{\Pi}_{\tilde{F}_{(b),(\text{con})}}$ is the projector onto the support of $\tilde{F}_{(b),(\text{con})}$. This projector plays a trivial rule in the measurement itself, and is only included to ensure that we obtain a valid POVM.
  3. 3.
    To obtain the common filtering operation, find a $\tilde{F}$ such that $\mathrm{I}\geq\tilde{F}\geq\tilde{F}_{(b_{A},b_{B}),(\text{con})}$ for all $(b_{A},b_{B})$. Then, set the second step basis-dependent filtering POVM $\{{F}_{(b),(\text{con})},{F}_{(b),(\bot)}\}$ to be
    ${F}_{(b),(\text{con})}$ $=\sqrt{\tilde{F}}^{+}\tilde{F}_{(b),(\text{con})}\sqrt{\tilde{F}}^{+}+\mathrm{I}-\mathrm{\Pi}_{\tilde{F}}\qquad\qquad(=\mathrm{I}-{F}_{(b),(\bot)})$ (76)
    ${F}_{(b),(\bot)}$ $=\sqrt{\tilde{F}}^{+}\left(\tilde{F}-\tilde{F}_{(b),(\text{con})}\right)\sqrt{\tilde{F}}^{+}$
  4. 4.
    Compute
    $\delta_{2}=$ $\norm{\mathrm{I}-{F}_{(Z),(\text{con})}}_{\infty}$
  5. 5.
    For the analysis of practical scenarios, where $\eta_{b_{i}},d_{b_{i}}$ are not known exactly but instead satisfy Eq. 74, one must also additionally maximize Eq. 77 over all possible choices of $\eta_{b_{i}},d_{b_{i}}$.
We construct the POVMs from Eqs. 75 and 76 in Section E.1. To bound $\delta_{1},\delta_{2}$, we use the fact that all POVMs are block-diagonal in the total photon number, and bound the $\infty$-norm of each block separately. Note that we can always treat the common value of loss in the detectors to be a part of the channel [52, Section III C]. This means that we pull out $(\max_{b,i}\{\eta_{b_{i}}\})$, and treat it as a part of the channel. This computation of $\delta_{1},\delta_{2}$ using the above steps is quite cumbersome, and is explained in Sections E.2 and E.3. Finally, we obtain
$\delta_{1}$ $\leq 4\max\left\{1-\sqrt{\frac{1-(1-d_{\text{min}})^{2}}{1-(1-d_{\text{max}})^{2}}},1-\sqrt{1-(1-d_{\text{min}})^{2}(1-r_{\eta})}\right\},$ (78)

### v2

@@ V.4 Bounds on zero and one photon statistics
The lower bound on the zero-photon component is given by [30, Eq. 2]
@@
The lower bound on the one-photon component is given by [30, Eq. 3]
@@
The upper bound on the one-photon component is given by [30, Eq. 4]
@@
Since $\Pr(\mathcal{S}_{\text{constraints}})\geq 1-9\varepsilon_{\text{AT-d}}^{2}$, and Eqs. 49, 50 and 51 follow from the expressions in $\mathcal{S}_{\text{constraints}}$. We have
@@
V.5 Variable-length security statement for decoy-state
Having proved Eq. 38, we now have the following theorem regarding variable-length security of the decoy-state BB84 protocol.
Theorem 3.
[ Variable-length security of decoy-state BB84] Suppose Eq. 38 is satisfied and let $\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ be a function that determines the number of bits used for error-correction in the QKD protocol. Define
$l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ :-$\coloneq\max\Bigg{(}0,\mathcal{B}_{1}\left(n_{K,\mu_{\vec{k}}}\right)\left(1-h\left(\mathcal{B}_{e}\left(e^{\text{obs}}_{X,\mu_{\vec{k}}},n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}}\right)\right)\right):$ (53)
$-\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})-2\log(1/2\varepsilon_{\text{PA}})-\log(2/\varepsilon_{\text{EV}})\Bigg{)}$
@@ V.5 Variable-length security statement for decoy-state
For instance, one way to avoid this problem is to argue that Bob can compare his raw key before and after error-correction to calculate the number of errors in the key generation rounds (assuming error-correction succeeded). This can indeed be made rigorous by arguing that if error-correction fails, the protocol aborts with high probability anyway (due to error-verification). However an additional issue remains. For variable-length protocols, Bob must announce either the number of errors he observes, or the length of key he wishes to produce, to Alice. This additional announcement leaks information to Eve which must be accounted for. Assuming that Bob announces the final output key length, a naive analysis would reduce the key length by an additional $\log(n_{\text{len}})$ where $n_{\text{len}}$ denotes the number of allowed output key length. These observations are missing in Ref. [50].
@@
Note that this problem is avoided by this work since the length of output key is a function only of the public announcements during Step 3 of the protocol (Section V.1).
VI Results
@@
We will now apply our results to a decoy-state BB84 protocol with realistic detectors. To do so, we start by outlining a recipe for using this work to compute key rates in Section VI.1. We will then specify the canonical model for our detectors with efficiency mismatch in Section VI.2. We will apply the recipe to our model in Section VI.3. Finally, we will plot the key rate we obtain in Section VI.4.
VI.1 Recipe for computing key rates in the presence of basis-efficiency mismatch
In this subsection, we provide straightforward instructions for using the results of this work to compute key rates for decoy-state BB84 in the presence of basis-efficiency mismatch (see the end of this subsection for a pointer to the exact expressions). We will start by explaining the computation of (upper bounds on) $\delta_{1},\delta_{2}$ for a given model of the measurement POVMs in the protocol. To do so, one has to break up the measurement process implemented by Alice and Bob into multiple steps via multiple uses of Lemma 1. This is done as follows:
  1. 1.
    Start with POVM $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ which describe Alice and Bob measuring in the $(b_{A},b_{B})$ basis, and obtaining a conclusive error, a conclusive no-error, and an inconclusive outcome respectively. (In this work, we apply this recipe on the POVMs defined in Eq. 58.)
  2. 2.
    Pick a $\tilde{F}\geq\Gamma_{(b_{A},b_{B}),(\neq)}+\Gamma_{(b_{A},b_{B}),(=)}$ for all $(b_{A},b_{B})$. Consider the four-outcome POVM
    $\{\mathrm{I}-\tilde{F},\tilde{F}-\Gamma_{(b_{A},b_{B}),(=)}-\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\neq)}\}$. Group the last three outcomes together, and use Lemma 1 to divide this measurement into two steps. In the first step, $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ is measured and latter outcomes discarded. The remaining rounds are measured using $\{\tilde{F}_{(b_{A},b_{B}),(\bot)},\tilde{F}_{(b_{A},b_{B}),(=)},\tilde{F}_{(b_{A},b_{B}),(\neq)}\}$ where
    $\tilde{F}_{(b_{A},b_{B}),(\bot)}$ $=\sqrt{\tilde{F}}^{+}(\tilde{F}-\Gamma_{(b_{A},b_{B}),(\neq)}-\Gamma_{(b_{A},b_{B}),(\neq)})\sqrt{\tilde{F}}^{+}+\mathrm{I}-\mathrm{\Pi}_{\tilde{F}}$ (54)
    $\tilde{F}_{(b_{A},b_{B}),(\neq)}$ $=\sqrt{\tilde{F}}^{+}\Gamma_{(b_{A},b_{B}),(\neq)}\sqrt{\tilde{F}}^{+}$
    $\tilde{F}_{(b_{A},b_{B}),(=)}$ $=\sqrt{\tilde{F}}^{+}\Gamma_{(b_{A},b_{B}),(=)}\sqrt{\tilde{F}}^{+}$
  3. 3.
    Consider the new POVM $\{\tilde{F}_{(b_{A},b_{B}),(\bot)},\tilde{F}_{(b_{A},b_{B}),(=)},\tilde{F}_{(b_{A},b_{B}),(\neq)}\}$ Using Lemma 1 again, divide this POVM measurement into two steps. The first step is implemented using $\{{F}_{(b_{A},b_{B}),(\text{con})},{F}_{(b_{A},b_{B}),(\bot)}\}$ and decides whether the outcome is conclusive or inconclusive. The conclusive outcomes are further measured using $\{G^{\text{con}}_{(b),(\neq)},G^{\text{con}}_{(b),(=)}\}$. These POVM elements are given by
    ${F}_{(b_{A},b_{B}),(\text{con})}$ $=\tilde{F}_{(b_{A},b_{B}),(\neq)}+\tilde{F}_{(b_{A},b_{B}),(=)}$ (55)
    ${F}_{(b_{A},b_{B}),(\bot)}$ $=\tilde{F}_{(b_{A},b_{B}),(\bot)}$
    $G^{\text{con}}_{(b_{A},b_{B}),(\neq)}$ $=\sqrt{{F}_{(b_{A},b_{B}),(\text{con})}}^{+}\tilde{F}_{(b_{A},b_{B}),(\neq)}\sqrt{{F}_{(b_{A},b_{B}),(\text{con})}}^{+}$
    $G^{\text{con}}_{(b_{A},b_{B}),(=)}$ $=\sqrt{{F}_{(b_{A},b_{B}),(\text{con})}}^{+}\tilde{F}_{(b_{A},b_{B}),(=)}\sqrt{{F}_{(b_{A},b_{B}),(\text{con})}}^{+}+\mathrm{I}-\mathrm{\Pi}_{{F}_{(b_{A},b_{B}),(\text{con})}}\qquad\qquad(=\mathrm{I}-G^{\text{con}}_{(b_{A},b_{B}),(\neq)})$
    where $\mathrm{\Pi}_{{F}_{(b_{A},b_{B}),(\text{con})}}$ is the projector onto the support of ${F}_{(b_{A},b_{B}),(\text{con})}$. This projector plays a trivial rule in the measurement itself, and is only included to ensure that we obtain a valid POVM.
  4. 4.
    Compute
    $\delta_{2}=$ $\norm{\mathrm{I}-{F}_{(Z),(\text{con})}}_{\infty}$
    where we recall that whenever the basis is explicitly written as $X/Z$, it represents both Alice and Bob’s basis choices.
  5. 5.
    For the analysis of practical scenarios, where $\eta_{b_{i}},d_{b_{i}}$ are not known exactly but are instead known to be in some range, one must also additionally maximize Eq. 56 over all possible choices of $\eta_{b_{i}},d_{b_{i}}$.
Once $\delta_{1},\delta_{2}$ are computed via the procedure above, we can compute key rates as follows. The key rate expression for the decoy-state BB84 protocol is given by Eq. 53. To use this expression, refer to Eqs. 41 and 38 (which are notationally equivalent). The bounds for the decoy analysis in Eq. 41 are in turn found in Eqs. 49, 50 and 51, whereas the bound for the phase error estimation is found in Eq. 34. For the BB84 protocol where Alice sends single photons, the key rate is given by Eqs. 6 and 34.
Note that the above recipe is valid for for all detector models, and yields non-trivial results as long as one can suitably bound $\delta_{1},\delta_{2}$. In this work, we consider the canonical model of detectors in the next section, and then compute the values of $\delta_{1},\delta_{2}$ for this model Section VI.3.
VI.2 Detector Model
@@
In this section, we specify the canonical model of Bob’s detectors (for active BB84) we use in this work. Let $\eta_{b_{i}},d_{b_{i}}$ denote the efficiency and dark count rate of Bob’s POVM corresponding to basis $b$, and bit $i$. We first define Bob’s double click POVM for basis $b\in\{Z,X\}$ to be $\Gamma^{(B)}_{(b,\text{dc})}=\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{0}})(1-\eta_{b_{0}})^{N_{0}})(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ , where |$\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ is the state with $N_{0}$ photons in the mode 1, and $N_{1}$ photons in mode 2, where the modes are defined with respect to basis $b$. For example, for polarisation-encoded BB84, |$\outerproduct{2,1}{2,1}_{Z}$ would signify the state with $2$ horizontally-polarised photons and $1$ vertically polarised photon. Recall that double clicks are mapped to single clicks randomly in our protocol. Thus, we can write Bob’s POVM elements as
@@
$\Gamma^{(B)}_{(b,1)}$ $=(1-d_{b_{0}})\sum_{N_{0},N_{1}=0}^{\infty}(1-\eta_{b_{0}})^{N_{0}}(1-(1-d_{b_{1}})(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}+\frac{1}{2}\Gamma^{(B)}_{(b,\text{dc})}.$
Decoy methods allow us to restrict out attention to rounds where Alice sent single photons. Thus her Hilbert space is qubit while Bob holds two optical modes. The joint Alice-Bob POVM elements for the basis $b$ can be constructed via Eqs. 2, 3 and 57 and are given by
@@
VI.3 Computing bounds on $\delta_{1},\delta_{2}$
In this subsection, we will compute upper bounds on $\delta_{1},\delta_{2}$ by following the recipe in Section VI.1.
VI.3.1 Active BB84 detection setup without any hardware modification
In this case the POVMs used by Alice and Bob are exactly given by Eq. 58. We construct the POVMs from Eqs. 55 and 54 in Section G.1. To bound $\delta_{1},\delta_{2}$, we use the fact that all POVMs are block-diagonal in the total photon number, and bound the $\infty$-norm of each block separately. Note that we can always treat the common value of loss in the detectors to be a part of the channel [52, Section III C]. This means that we pull out $(\max_{b,i}\{\eta_{b_{i}}\})$, and treat it as a part of the channel. (This is equivalent to giving the $\{\tilde{F},\mathrm{I}-\tilde{F}\}$ measurement to Eve.) This computation of $\delta_{1},\delta_{2}$ using the above steps is quite cumbersome, and is explained in Sections G.2 and G.3. Finally, we obtain
$\delta_{1}$ $\leq\max\left\{\left(1-\frac{1-(1-d_{\text{min}})^{2}}{1-(1-d_{\text{max}})^{2}}\right)\frac{d_{\text{max}}(2-d_{\text{min}})}{1-(1-d_{\text{min}})^{2}},4\absolutevalue{1-\sqrt{1-(1-d_{\text{min}})^{2}(1-r_{\eta})}}\right\},$ (60)